OpenStudy (anonymous):

Gilda walks to the train station. If she walks at the rate of 3 mph, she misses her train by 7 minutes. However, if she walks at the rate of 4 mph, she reaches the station 5 minutes before the arrival of the train. Find the distance Gilda walks to the station. David knew he made a mistake when he calculated that Gilda walks 123 miles to the station. David's calculations: Using d = rt, the distance is the same, but the rate and time are different. If Gilda misses the train, it means the time t needs 7 more minutes so d = 3(t + 7). If she gets to the station 5 minutes early

OpenStudy (anonymous):

means the time t can be 5 minutes less so d = 4(t - 5). 3(t + 7) = 4(t - 5) 3t + 21 = 4t - 20 t = 41 d = rt, so d = 3(41) = 123 Find David's mistake in his calculations. In two or more complete sentences, explain his mistake. Include the correct calculations and solutions in your answer.